# Factoriality of groupoid von Neumann algebras

### 06 Feb 2024 v1

@@ Abstract
We give a characterisation of factoriality of the groupoid von Neumann algebra $L(\mathcal{G})$ associated to a discrete measured groupoid $(\mathcal{G},\mu)$. We introduce the notion of groupoids with ‘infinite conjugacy classes’ and show that this property together with ergodicity of the groupoid is equivalent to factoriality of $L(\mathcal{G})$.
@@ 1 Introduction
The organisation of the first part of the paper is as follows. We state some preliminaries on discrete measured groupoids and their von Neumann algebras in Section 2, and then gather some folklore results on bases of groupoids in Section 3. A lot of these results have already been used intrinsically in the literature, for example in [Ana13]. In Section 4, we define the notion of conjugacy classes for Borel subsets of the isotropy subgroupoid and we introduce an infinite conjugacy class (icc) condition for groupoids (see Definition 4.2). Then we prove the following theorem, which is the main result of this paper.
@@
Similarly for factoriality, beyond the principal case the 2-cocycle creates a non-trivial obstruction. A characterization of factoriality for twisted group von Neumann algebras was given by Kleppner in [Kle62]. In Section 5.3, we introduce the notion of central subsets of the isotropy subgroupoid and in Theorem 5.5, we give a characterization of factoriality for cocycle twisted groupoid von Neumann algebras in terms of these central subsets. We also introduce a similar condition as in [Kle62] for twisted discrete measure groupoids. In Lemma 5.8 we show that the von Neumann algebra associated to a pair $(\mathcal{G},\omega)$ consisting of an ergodic groupoid and a 2-cocycle that satisfies Kleppner’s condition must indeed be a factor.
@@ 3 Basis of a groupoid, Fourier decomposition, and the $j$ map
$\bigcup_{B}CBC^{-1}=C\left(\bigcup_{B}B\right)C^{-1}=C\mathcal{G}C^{-1}\;.$
@@ 4 Factoriality of the untwisted groupoid von Neumann algebra
We say $(\mathcal{G},\mu)$ has $\mu$-infinite conjugacy classes, or simply that it is icc, if whenever $A\subset\operatorname{Iso}(\mathcal{G})$ is a bisection (or equivalently, whenever $A$ is a Borel subset) with finite measure conjugacy class, then $A\subset\mathcal{G}^{(0)}$.
@@
If a Borel set $A$ has a $\mu$-finite conjugacy class $\Omega_{A}$, one may ask if, like the case of groups, $\Omega_{A}$ is the disjoint union of finitely many conjugates of $A$. Although the general case is still open, if the groupoid is assumed to be ergodic, the following lemma shows that $\Omega_{A}$ is the union of finitely many mutually disjoint global bisections. In particular, in that case we have that $\mu_{s}(\Omega_{A})=\mu_{t}(\Omega_{A})$.
@@
Let $(\mathcal{G},\mu)$ be an ergodic measured groupoid and let $A\subset\operatorname{Iso}(\mathcal{G})$ be a Borel subset with $\mu$-finite conjugacy class. Then there exist finitely many mutually disjoint global bisections $V_{1},...,V_{k}$ such that $\Omega_{A}=\bigsqcup_{i}V_{i}$. In particular, $\mu_{s}(\Omega_{A})=\mu_{t}(\Omega_{A})=k$.
@@
By ergodicty we get that for each $k$ either $\mu(E_{k})=0$ or $\mu(E_{k})=1$. Since $\mathcal{G}^{(0)}=\bigsqcup_{k}E_{k}$, we find a $k$ for which $\mu(E_{k})=1$. Then, by repeatedly applying the Lusin–Novikov theorem we find global bisections $V_{1},...,V_{k}\subset\operatorname{Iso}(\mathcal{G})$ such that $\Omega_{A}=\bigsqcup_{i}V_{i}$. ∎
@@
$\operatorname{supp}(j(a))=\bigcup_{n\geq 1}\left\{g\in\mathcal{G}\ \left|\ |j(a)|^{2}\geq\frac{1}{n^{2}}\right.\right\},$
and fix an $n\in\mathbb{N}$ such that $E:=\left\{g\in\mathcal{G}\ |\ |j(a)|^{2}\geq\frac{1}{n^{2}}\right\}$ satisfies $\mu_{s}(E)>0$. Note that $\mu_{s}(E\cap\mathcal{G}^{(0)})=0$. By Lemma 3.10 we know that $E\subset\operatorname{Iso}(\mathcal{G})$ and from Lemma 3.11 it follows that $g\cdot E\cdot g^{-1}\subset E$ for all $g\in\mathcal{G}$. Finally, to see that $\mu_{s}(E)<\infty$, note that
@@ 6.1 Group bundles
We will prove that $\Gamma_{x}$ is icc for a.e. $x\in X$ if and only if every nontrivial bisection $B\subset\mathcal{G}$ has $\mu$-infinite conjugacy class. Then the result follows from Theorem A. For an element $g\in\mathcal{G}$ we write $C_{g}$ for the conjugacy class of $g$ in $\Gamma_{s(g)}$. Write $Y=\{x\in\mathcal{G}^{(0)}:\Gamma_{x}\text{ is not icc}\}$, which is Borel by Lemma 6.2 below.
First assume that $\mu(Y)>0$. By Lemma 6.2 below there exists a section $\sigma\colon\mathcal{G}^{(0)}\rightarrow\mathcal{G}$ such that for every $x\in Y$ we have that $\sigma(x)\neq e_{x}$ and $\sigma(x)$ has finite conjugacy class. Define $Y_{n}=\{x\in Y:|C_{\sigma(x)}|=n\}$, so that $Y=\bigsqcup_{n}Y_{n}$. Then there exists an $n\in\mathbb{N}$ such that $\mu(Y_{n})>0$. Now define the section $\tau\colon\mathcal{G}^{(0)}\rightarrow\mathcal{G}$ by
$\tau(x)=\begin{cases}\sigma(x)&\text{ if }x\in Y_{n},\\ e_{x}&\text{ if }x\notin Y_{n}.\end{cases}$ (17)
By construction the set $\bigcup_{g\in\mathcal{G}}g\tau(\mathcal{G}^{(0)})g^{-1}$ has measure $\mu(\mathcal{G}^{(0)}\setminus Y_{n})+n\mu(Y_{n})<+\infty$. So $\tau(\mathcal{G}^{(0)})$ is a a nontrivial bisection with $\mu$-finite conjugacy class.
Conversely, assume that $B\subset\mathcal{G}$ is a nontrivial bisection with $\mu$-finite conjugacy class. For $x\in s(B)$ write $b_{x}$ for the unique element $b\in B$ with $s(b)=x$. We have that
$\bigcup_{g\in\mathcal{G}}gBg^{-1}=\bigsqcup_{x\in s(B)}C_{b_{x}}\,.$ (18)
So, by assumption, we get that
[…]
In particular we have that $|C_{b_{x}}|<+\infty$ for a.e. $x\in s(B)$. Since $B$ is nontrivial, this implies that $\mu(Y)>0$. ∎
@@ 6.2 Transformation groupoids
Conversely, assume that $G\curvearrowright X$ is induced from $H\curvearrowright Y$ and that $h\in H\setminus\{h\}$ acts trivially on $Y$ such that $h$ has finite conjugacy class $C$. Then we have that
@@
So $\{h\}\times Y$ is a nontrivial partial bisection that has $\mu$-finite conjugacy class, contradicting statement $2$ of the proposition. ∎
@@ 6.3 Full subgroupoids and partial actions
If $G\curvearrowright X$ is induced from $H\curvearrowright Y$, then $Y$ is full by definition. Conversely, note that by fullness of $Y$, that $\bigcup_{g\in G}gY=t((G\ltimes X)Y)=X$. To show these sets are mutually disjoint, let $g\not\in H$. If there is some $y\in gY\cap Y$, then $(g,y)\in(G\ltimes X)|_{Y}=H\ltimes Y$. Thus, $g\in H$, reaching a contradiction. The measurable setting is the same, with equality replaced with equality up to a negligible subset. ∎
@@ 6.4 Deaconu–Renault Groupoids
$s(\bigcup_{n\neq 0}B_{n})=\{x\;|\;\text{there exists }k>l\geq 0\text{ such that }\sigma^{k}(x)\neq\sigma^{l}(x)\}\;.$

### 27 Feb 2024 v2

@@ Abstract
We give a characterisation of the factoriality of the groupoid von Neumann algebra $L(\mathcal{G})$ associated to a discrete measured groupoid $(\mathcal{G},\mu)$. We introduce the notion of groupoids with ‘infinite conjugacy classes’ and show that this property together with ergodicity of the groupoid is equivalent with factoriality of $L(\mathcal{G})$.
@@ 1 Introduction
The organisation of the first part of the paper is as follows. We state some preliminaries on discrete measured groupoids and their von Neumann algebras in Section 2, and then gather some folklore results on bases of groupoids in Section 3. A lot of these results have already been used intrinsically in the literature, for example in [Ana13]. In Section 4, we define the notion of conjugacy classes for Borel subsets of the isotropy subgroupoid. For such a Borel subset $E$ we show that the set $\Omega_{E}$ consisting of all conjugates of $E$ is a Borel subset and we call such a conjugacy class finite if it has finite measure. In Definition 4.2 we define groupoids satisfying the infinite conjugacy class (icc) condition if any such subset with a finite conjugacy class is essentially a subset of the unit space. Then we prove the following theorem, which is the main result of this paper.
@@
Similarly for factoriality, beyond the principal case the 2-cocycle creates a non-trivial obstruction. A characterization of factoriality for twisted group von Neumann algebras was given in [Kle62]. In Section 5.3, we introduce the notion of central subsets of the isotropy subgroupoid. If $\omega$ is a 2-cocycle on the groupoid, we call such a conjugation invariant Borel subset $E\subset\operatorname{Iso}(\mathcal{G})$ central if there exists a function $f:E\rightarrow\mathbb{C}$ such that:
$f(ghg^{-1})=\overline{\omega(ghg^{-1},g)}\omega(g,h)f(h)$
for all $h\in E$ and all elements $g$ with the same source as $h$. In Theorem 5.5 and Corollary 5.6, we give a characterization of factoriality for cocycle twisted groupoid von Neumann algebras in terms of these central subsets. We also introduce a similar condition as in [Kle62] for twisted discrete measure groupoids, and show that it is a necessary condition for the factoriality of the twisted groupoid von Neumann algebra.
@@
Acknowledgements: The authors would like to thank Stefaan Vaes for his valuable comments on the article.
@@ 3 Basis of a groupoid, Fourier decomposition, and the $j$ map
$\bigcup_{B}CBC^{-1}=C\left(\bigcup_{B}B\right)C^{-1}=C\mathcal{G}C^{-1}\;.\qed$
@@ 4 Factoriality of the untwisted groupoid von Neumann algebra
We say $(\mathcal{G},\mu)$ has $\mu$-infinite conjugacy classes, or simply that it is icc, if whenever $A\subset\operatorname{Iso}(\mathcal{G})$ is a nontrivial bisection (or equivalently, whenever $A$ is a non-null Borel subset) with finite measure conjugacy class, then $A\subset\mathcal{G}^{(0)}$.
@@
Since the source and target map coincide on $\operatorname{Iso}(\mathcal{G})$ we get that $\mu_{s}(\Omega_{A})=\mu_{t}(\Omega_{A})$ for every bisection $A\subset\operatorname{Iso}(\mathcal{G})$. In partcicular, it does not matter if we use the measure $\mu_{s}$ or $\mu_{t}$ in Definition 4.2 above.
@@
Let $(\mathcal{G},\mu)$ be an ergodic measured groupoid and let $A\subset\operatorname{Iso}(\mathcal{G})$ be a Borel subset with $\mu$-finite conjugacy class. Then there exist $k\in\mathbb{N}$ and finitely many mutually disjoint global bisections $V_{1},...,V_{k}$ such that $\Omega_{A}=\bigsqcup_{i}V_{i}\mod\mu_{s}$. In particular, $\mu_{s}(\Omega_{A})=k$.
@@
By ergodicty we get that for each $k$ either $\mu(E_{k})=0$ or $\mu(E_{k})=1$. Since $\mathcal{G}^{(0)}=\bigsqcup_{k}E_{k}$, we find a $k$ for which $\mu(E_{k})=1$. Then, by repeatedly applying the Lusin–Novikov theorem we find global bisections $V_{1},...,V_{k}\subset\operatorname{Iso}(\mathcal{G})$ such that $\Omega_{A}=\bigsqcup_{i}V_{i}\mod\mu_{s}$. ∎
Remark 4.5.
Removing the ergodicity assumption in Lemma 4.4, it is straightforward to come up with a counterexample. The transposition $(1,2)$ has conjugacy class of size (${n\choose 2}$ in the symmetric group $\mathcal{S}_{n}$. Then writing $\rho=\sum_{n=2}^{\infty}2^{-n}{n\choose 2}^{-1}<+\infty$ we get that $\mu(n)=\rho^{-1}2^{-n}{n\choose 2}^{-1}$ defines a probability measure $\mu$ on $\mathbb{N}_{\geq 2}$. Define $\mathcal{G}=\{(n,\sigma)\mid n\geq 2\text{ and }\sigma\in\mathcal{S}_{n}\}$ with unit space $(\mathbb{N}_{\geq 2},\mu)$ and source and target map given by $(n,\sigma)\mapsto n$. The conjugacy class $\Omega_{A}$ of the bisection $A=\bigcup_{n\geq 2}\{(n,(1,2))\}$ satisfies $\mu(\Omega_{A})<+\infty$ by construction, but it cannot be written as the union of finitely many bisections.
It immediately follows from the proof of Lemma 4.4 above that an ergodic discrete measured groupoid $\mathcal{G}$ is icc if and only if for every bisection $A\subset\operatorname{Iso}(\mathcal{G})$ we have that $|s^{-1}(x)\cap\Omega_{A}|=+\infty$ for a.e. $x\in\mathcal{G}^{(0)}$, thus depending only on the measure class of $\mu$. In general we have the following characterization that is independent of the choice of equivalent measure.
Proposition 4.6.
Suppose that $(\mathcal{G},\mu)$ is a discrete measured groupoid. For a bisection $A\subset\operatorname{Iso}(\mathcal{G})$ and an element $x\in\mathcal{G}^{(0)}$ write $A_{x}=\bigcup_{g\in t^{-1}(x)}gAg^{-1}=s^{-1}(x)\cap\Omega_{A}$. The following are equivalent:
  1. 1.
    The groupoid $(\mathcal{G},\mu)$ is icc.
  2. 2.
    For every bisection $A\subset\operatorname{Iso}(\mathcal{G})\setminus\mathcal{G}^{(0)}$ we have that $|A_{x}|=+\infty$ for a.e. $x\in s(A)$.
Proof.
First assume that every bisection $A\subset\operatorname{Iso}(\mathcal{G})\setminus\mathcal{G}^{(0)}$ satisfies $|A_{x}|=+\infty$ for a.e. $x\in s(A)$. Then for every bisection $A\subset\operatorname{Iso}(\mathcal{G})\setminus\mathcal{G}^{(0)}$ with $\mu_{s}(A)>0$ we get that
[…]
so that $\mathcal{G}$ is icc.
Conversely, assume that $A\subset\operatorname{Iso}(\mathcal{G})\setminus\mathcal{G}^{(0)}$ is a bisection such that
$\mu(\{x\in s(A)\mid|A_{x}|<+\infty\})>0\,.$
Then there exists $N\in\mathbb{N}$ large enough such that the set $B=\{g\in A\mid|A_{s(g)}|\leq N\}$ satisfies $\mu_{s}(B)>0$. Because $B\subset A$ we get that $|B_{x}|\leq|A_{x}|$ for every $x\in s(B)$. Furthermore, since $B_{k\cdot x}=kB_{x}k^{-1}$ we get that $|B_{x}|\leq N$ for every $x\in s(\Omega_{B})$. Then we see that
[…]
showing that $\mathcal{G}$ is not icc. ∎
@@
$\operatorname{supp}(j(a))=\bigcup_{n\geq 1}\left\{g\in\mathcal{G}\ \left|\ |j(a)(g)|^{2}\geq\frac{1}{n^{2}}\right.\right\},$
and fix an $n\in\mathbb{N}$ such that $E:=\left\{g\in\mathcal{G}\ |\ |j(a)(g)|^{2}\geq\frac{1}{n^{2}}\right\}$ satisfies $\mu_{s}(E)>0$. Note that $\mu_{s}(E\cap\mathcal{G}^{(0)})=0$. By Lemma 3.10 we know that $E\subset\operatorname{Iso}(\mathcal{G})$ and from Lemma 3.11 it follows that $g\cdot E\cdot g^{-1}\subset E$ for all $g\in\mathcal{G}$. Finally, to see that $\mu_{s}(E)<\infty$, note that
@@ 6.1 Group bundles
Write $Y=\{x\in\mathcal{G}^{(0)}:\Gamma_{x}\text{ is not icc}\}$, which is Borel by Lemma 6.2 below. For an element $g\in\mathcal{G}$ we write $C_{g}$ for the conjugacy class of $g$ in $\Gamma_{s(g)}$. We will prove that $\mu(Y)=0$ if and only if for every nontrivial bisection $A\subset\mathcal{G}\setminus\mathcal{G}^{(0)}$ we have that $|C_{g}|=+\infty$ for a.e. $g\in A$. Then the result follows from Proposition 4.6 and Theorem A.
First assume that $\mu(Y)>0$. By Lemma 6.2 below there exists a section $\sigma\colon\mathcal{G}^{(0)}\rightarrow\mathcal{G}$ such that for every $x\in Y$ we have that $\sigma(x)\neq e_{x}$ and $\sigma(x)$ has finite conjugacy class. Then $\sigma(Y)\subset\mathcal{G}\setminus\mathcal{G}^{(0)}$ is a nontrivial bisection and $|C_{g}|<+\infty$ for every $g\in\sigma(Y)$ by construction.
Conversely, assume that $\mu(Y)=0$ and let $A\subset\mathcal{G}\setminus\mathcal{G}^{(0)}$ be a nontrivial bisection. By definition $|C_{g}|=+\infty$ for every $g\in A\cap s^{-1}(\mathcal{G}^{(0)}\setminus Y)$, which is a conull set inside $A$. ∎
@@ 6.2 Transformation groupoids
Conversely, assume that $G\curvearrowright X$ is induced from $H\curvearrowright Y$ and that $h\in H\setminus\{e\}$ acts trivially on $Y$ such that $h$ has finite conjugacy class $C$. Then we have that
@@
So $\{h\}\times Y$ is a partial bisection that has $\mu$-finite conjugacy class, contradicting statement $2$ of the proposition. ∎
@@ 6.3 Full subgroupoids and partial actions
If $G\curvearrowright X$ is induced from $H\curvearrowright Y$, then $Y$ is full by definition. Conversely, note that by fullness of $Y$, that $\bigcup_{g\in G}gY=t((G\ltimes X)Y)=X$. To show these sets are mutually disjoint, let $g\not\in H$. If there is some $y\in gY\cap Y$, then $(g,y)\in(G\ltimes X)|_{Y}=H\ltimes Y$. Thus, $g\in H$, reaching a contradiction. The measurable setting is the same, with equality replaced with equality up to a negligable subset. ∎
@@ 6.4 Deaconu–Renault Groupoids
$s(\bigcup_{n\neq 0}B_{n})=\{x\;|\;\text{there exists }k>l\geq 0\text{ such that }\sigma^{k}(x)\neq\sigma^{l}(x)\}\;.\qed$